OpenStudy (anonymous):

2 microF capacitor is charged to 50V and then connected in parallel (positive to positive plate) with a 4 microF capacitor charged to 100 V. what are the final charges on the capacitors? what is the potential difference across each?

OpenStudy (anonymous):

the answers are: (a.) 167microF , 333 microF (b.) 83 V. how can i arrive these answers?

OpenStudy (anonymous):

First of all, you should find total charge by using Q=C*V Voltage is same both end of the capacitors, because they are connected as parallel. V1=V2 \[\frac{ Q1 }{ 2muf } = \frac{ Q2 }{ 4muf }\] And you know the total charged, Qtotal=Q1+Q2 then you can calculate Q1 and Q2. For calculate the voltage, you can use first equation again.

OpenStudy (anonymous):

in finding the total charge, what voltage will i use? the first or the second one?

OpenStudy (anonymous):

it does not matter. When two capacitors are connected as parallel, Voltages are same both capacitors' end.

OpenStudy (anonymous):

Initial Energy = Final Energy

OpenStudy (anonymous):

1/2 C1 V1^2 + 1/2 C2 V2^2 = 1/2 Cnet * V^2